# A Block-Shifted Cyclic Reduction Algorithm for Solving a Class of Quadratic Matrix Equations

### v1

@@ 5 Combination of CR, block shift-and-deflate, and QZ algorithm
Under our assumptions, the matrix polynomial $B(z)=B_{0}+B_{1}z+B_{2}z^{2}$ associated with (19) has eigenvalues $\mu_{1},\ldots,\mu_{\ell}$, each with multiplicity 2, and $\bar{G}_{11}$ is the only solution to (19) with eigenvalues $\mu_{1},\ldots,\mu_{\ell}$. To compute $\bar{G}_{11}$, by borrowing the idea from [11], we employ the generalized Schur decomposition, which can be computed via the QZ algorithm [12, Chapter 7], and derive the following Algorithm 2. Similarly, $\bar{R}_{11}$ is the only solution to (22) with eigenvalues $1/\mu_{1},\ldots,1/\mu_{\ell}$. Analogous to (8), $\bar{R}_{11}$ can be computed from
@@
Solve the $\ell\times\ell$ quadratic matrix equations(19) by means of the Algorithm 2 and obtain $\bar{G}_{11}$.
@@ 6 Numerical results
We compare the proposed Algorithm 3, denoted by BS-CR, with standard CR, denoted by CR and described in [3, Section 7.5], the Shifted CR, denoted by S-CR and proposed in [30], and the $U$-based fixed point iteration FPI with starting approximation a stochastic matrix (see [3, Chapter 6]).
@@ 6.2 Example 2
$E_{0}=\begin{bmatrix}0&S_{1}\\ S_{2}&0\end{bmatrix},\penalty 10000\ \penalty 10000\ E_{1}=0,\penalty 10000\ \penalty 10000\ E_{2}=\begin{bmatrix}0&S_{2}\\ S_{1}&0\end{bmatrix}$,
where $S_{1}\geq 0$ and $S_{2}\geq 0$ are $p\times p$ matrices such that $(S_{1}+S_{1})\mathbf{1}=\mathbf{1}$. This example generalizes the example on page 133 in [35]. Here we take
$S_{1}=\frac{1}{8}\begin{bmatrix}3&1\\ 1&2&1\\ &\ddots&\ddots&\ddots\\ &&1&2&1\\ &&&1&3\end{bmatrix},\penalty 10000\ \penalty 10000\ S_{2}=\frac{1}{10}\begin{bmatrix}4&1\\ 1&3&1\\ &\ddots&\ddots&\ddots\\ &&1&3&1\\ &&&1&4\end{bmatrix}$.
The matrix polynomial $A(z)$ has $\ell=2$ double eigenvalues at the square root of $1$. We test the robustness of different versions of the CR algorithm (BS-CR, S-CR, and CR) for increasing values of the size $p$. In our implementations, the iterations are terminated once the maximum number of iterations is reached, with $k_{\max}=100$ for CR and S-CR, and $k_{\max}=12$ for BS-CR. The numerical results are presented in Figure 1, which displays the residual errors for the three algorithms. As observed from the figure, S-CR for some case provides an error of the order $10^{0}$, CR provides approximation of approximately of the order of the square root of the machine precision, while BS-CR achieves the highest accuracy, of the order $10^{-15}$.
@@
In Figure 2 we report the CPU time and the number of iterations needed by the different versions of CR. As we expect, S-CR and CR require a lower CPU time with respect to BS-CR. However, the asymptotic growth of the 3 algorithms almost the same. Indeed, the larger computational cost of each step of BS-CR is balanced by the lower number of iterations; moreover CR always reached the maximum number of iteration and, even performing more iterations, the residual does not improve.
@@ 6.3 Example 3
$A_{0}=PG,\penalty 10000\ \penalty 10000\ A_{1}=-RPG-P\quad\text{and}\quad A_{2}=RP,$
where $P$ is a nonsingular matrix. By construction, $G$ and $R$ are solutions of the QMEs (3) and (4), respectively. We choose
$P=\operatorname{tridiag}(-1,4,-1)\in\mathbb{R}^{m\times m},$
@@ Acknowledgements
The work on this paper was done during the first author’s visit to Prof. Beatrice Meini at the University of Pisa. The first author is deeply grateful to Prof. Meini for her hospitality and guidance.
@@
    The contribution of the authors is parithetic.

### v2

@@ 5 Combination of CR, block shift-and-deflate, and QZ algorithm
Under our assumptions, the matrix polynomial $B(z)=B_{0}+B_{1}z+B_{2}z^{2}$ associated with (19) has eigenvalues $\mu_{1},\ldots,\mu_{\ell}$, each with multiplicity 2, and $\bar{G}_{11}$ is the only solution to (19) with eigenvalues $\mu_{1},\ldots,\mu_{\ell}$. To compute $\bar{G}_{11}$, by borrowing the idea from [21], we employ the generalized Schur decomposition, which can be computed via the QZ algorithm [15, Chapter 7], and derive the following Algorithm 2. Similarly, $\bar{R}_{11}$ is the only solution to (22) with eigenvalues $1/\mu_{1},\ldots,1/\mu_{\ell}$. Analogous to (8), $\bar{R}_{11}$ can be computed from
@@
Solve the $\ell\times\ell$ quadratic matrix equation (19) by means of the Algorithm 2 and obtain $\bar{G}_{11}$.
@@ 6 Numerical results
We compare the proposed Algorithm 3, denoted by BS-CR, with standard CR, denoted by CR and described in [7, Section 7.5], the Shifted CR, denoted by S-CR and proposed in [20], and the $U$-based fixed point iteration FPI with starting approximation a stochastic matrix (see [7, Chapter 6]).
@@ 6.2 Example 2
$E_{0}=\begin{bmatrix}0&F_{1}\\ F_{2}&0\end{bmatrix},\penalty 10000\ \penalty 10000\ E_{1}=0,\penalty 10000\ \penalty 10000\ E_{2}=\begin{bmatrix}0&F_{2}\\ F_{1}&0\end{bmatrix}$,
where $F_{1}\geq 0$ and $F_{2}\geq 0$ are $p\times p$ matrices such that $(F_{1}+F_{2})\mathbf{1}=\mathbf{1}$. This example generalizes the example on page 133 in [14]. Here we take
$F_{1}=\frac{1}{8}\begin{bmatrix}3&1\\ 1&2&1\\ &\ddots&\ddots&\ddots\\ &&1&2&1\\ &&&1&3\end{bmatrix},\penalty 10000\ \penalty 10000\ F_{2}=\frac{1}{10}\begin{bmatrix}4&1\\ 1&3&1\\ &\ddots&\ddots&\ddots\\ &&1&3&1\\ &&&1&4\end{bmatrix}$.
The matrix polynomial $A(z)$ has $\ell=2$ double eigenvalues at the square root of $1$. We test the robustness of different versions of the CR algorithm (BS-CR, S-CR, and CR) for increasing values of the size $p$. In our implementations, the iterations are terminated once the maximum number of iterations is reached, with $k_{\max}=100$ for CR and S-CR, and $k_{\max}=12$ for BS-CR. The numerical results are presented in Figure 1, which displays the residual errors for the three algorithms. As observed from the figure, S-CR for some case provides an error of the order $10^{0}$, CR produces an approximation of the order of the square root of the machine precision, while BS-CR achieves the highest accuracy, of the order $10^{-15}$.
@@
In Figure 2 we report the CPU time and the number of iterations needed by the different versions of CR. As we expect, S-CR and CR require a lower CPU time with respect to BS-CR. However, the asymptotic growth of the three algorithms almost the same. Indeed, the larger computational cost of each step of BS-CR is balanced by the lower number of iterations; moreover CR always reached the maximum number of iteration and, even performing more iterations, the residual does not improve.
@@ 6.3 Example 3
$A_{0}=HG,\penalty 10000\ \penalty 10000\ A_{1}=-RHG-P\quad\text{and}\quad A_{2}=RH,$
where $H$ is a nonsingular matrix. By construction, $G$ and $R$ are solutions of the QMEs (3) and (4), respectively. We choose
$H=\operatorname{tridiag}(-1,4,-1)\in\mathbb{R}^{m\times m},$
@@ Acknowledgements
The authors appreciate the constructive suggestions provided by the referees, which have significantly enhanced the clarity and quality of the paper. This work was initiated and primarily conducted during the first author’s visit to the University of Pisa. The first author is deeply grateful to Prof. Beatrice Meini for her kind invitation and hospitality, and to the Department of Mathematics for providing the excellent research environment that made this collaboration possible.
@@
    The contributions of the authors are equal.